Two pipes together can fill a large tank in 10 hours. One of the pipes, used alone, takes 15 hours longer than the other to fill the tank. How long would each pipe used alone take to fill the tank
Answer by checkley77(12569)

(15+X)X/(15+X+X)=10
(15X+X^2)/(15+2X)=10
15X+X^2=10(15+2X)
X^2+15X=150+20X
X^2+15X-20X-150=0
X^2-5X-150=0
(X-15)(X+10)=0
X=15 HOURS FOR THE FIRST PIPE.
15+15=30 HOURS FOR THE SECOND PIPE TO FILL THE TANK.
PROOF:
30*15/(30+15)10
450/45=10
10=10